# Testing New Physics in Oscillations at a Neutrino Factory

### v1

@@ 5 Conclusion
We find that combining 10 years of DUNE with 10 years of a neutrino factory will improve over the constraints on these new physics scenarios from DUNE alone even when considering 20 years of DUNE. Furthermore, the addition of a neutrino factory can remove degeneracies in the DUNE-only constraints due to the presence of new oscillation channels, especially the $\nu_{e}$ disappearance and $\nu_{\mu}$ appearance channels.
@@ Acknowledgements.
PBD acknowledges support by the United States Department of Energy under Grant Contract No. DE-SC0012704. JG acknowledges support by the U.S. Department of Energy Office of Science under award number DE-SC0025448. CFK acknowledges the usage of INPAC cluster at Shanghai Jiao Tong University.
@@ Appendix A Additional NSI results
In Fig. 11 we show the precision on $\delta_{CP}$ assuming $\delta_{CP}^{\rm{true}}=0$ in the presence of NSI. Figs. 12, 13, 14, 15 we show the sensitivity to $\theta_{23}$ in the presence of non-zero NSI parameters, similar to Fig. 6 in the main text.

### v2

@@ 5 Conclusion
We find that combining 10 years of DUNE with 10 years of a neutrino factory will improve over the constraints on these new physics scenarios from DUNE alone even when considering 20 years of DUNE. Furthermore, the addition of a neutrino factory can remove numerous degeneracies in the DUNE-only constraints due to the presence of new oscillation channels, especially the $\nu_{e}$ disappearance and $\nu_{\mu}$ appearance channels.
@@ Acknowledgements.
PBD acknowledges support by the United States Department of Energy under Grant Contract No. DE-SC0012704. JG acknowledges support by the U.S. Department of Energy Office of Science under award number DE-SC0025448. CFK acknowledges support by the National Natural Science Foundation of China (12425506, 12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604). CFK also acknowledges the usage of INPAC cluster at Shanghai Jiao Tong University.
@@ Appendix A Additional NSI results
In this appendix we present additional results. Figs. 11, 12, 13, 14, 15, 16, 17, 19 show the results for the BNL setup, similar to figs. 1, 2, 3, 5, 7, 9, 10, 6 in the main text. In Fig. 18 we show the precision on $\delta_{CP}$ assuming $\delta_{CP}^{\rm{true}}=0$ in the presence of NSI. Figs. 20, 21, 22, 23 we show the sensitivity to $\theta_{23}$ in the presence of non-zero NSI parameters, similar to Fig. 6 in the main text.
Figure 11: Same as Fig. 1 but for the BNL configuration.
Figure 12: Same as Fig. 2 but for the BNL configuration.
Figure 13: Same as Fig. 3 but for the BNL configuration.
Figure 14: Same as Fig. 5 but for the BNL configuration.
Figure 15: Same as Fig. 7 but for the BNL configuration.
Figure 16: Same as Fig. 9 but for the BNL configuration.
Figure 17: Same as Fig. 10 but for the BNL configuration.